OpenStudy (anonymous):

How would I go about solving for the sum of telescoping series : ln 1/2 + ln 2/3 + ln 3/4

OpenStudy (anonymous):

ln1 - ln2 + ln2 - ln3 + ln3 - ln4 = ln1 - ln4 = ln(1/4)